OpenStudy (anonymous):

how much louder is 103dB than 100dB?

OpenStudy (anonymous):

okay, so I know that to produce a sound that is twice as loud we need a wave that is about ten times the intensity.

OpenStudy (anonymous):

really, i didn't know that.

OpenStudy (anonymous):

so 103dB is 3 times as loud?

OpenStudy (anonymous):

lol yes, really.. wait.. so maybe i should find the intensity of each and compare?

OpenStudy (anonymous):

yes. that's waht i was thinking.

OpenStudy (anonymous):

*what.

OpenStudy (anonymous):

brb.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think is is approx 10 times louder.

OpenStudy (anonymous):

they said it should be 2 times greater

OpenStudy (anonymous):

i used IL=10log(I/I0) & L=k log I to find that out. What do u think?

OpenStudy (anonymous):

Then, i have no idea how this works. :\

OpenStudy (anonymous):

i used I=(1x10^-12)(10^10.3) and I2=(1x10^-12)(10^10) then divided..

OpenStudy (anonymous):

okay, then we'll just leave that one out.

OpenStudy (anonymous):

Ahhh.. i think ur approach is better.

OpenStudy (anonymous):

okay, i don't have my calculator.. if you work out the result.. i got 0.3

OpenStudy (anonymous):

Heyyy, i got 2 when i divided them as u said!!!

OpenStudy (anonymous):

really! how do you divide without a calulator?

OpenStudy (anonymous):

<< i did use a calculator..>>

OpenStudy (anonymous):

lol but if we didn't.. lemme right them better

OpenStudy (anonymous):

\[I _{1}= (1x 10^{-12})(10^{10.3}) = \]

OpenStudy (anonymous):

do i just add the powers?

OpenStudy (anonymous):

No, when u divide I1 by I2, 10^-12 cancels out. So, keep it as it is.

OpenStudy (anonymous):

ahh, so then we just subtract the powers with the tens?\[\frac{ 10^{10.3} }{ 10^{^10}}\]

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

so 1^.3?

OpenStudy (anonymous):

10^0.3

OpenStudy (anonymous):

ahh!! used the calculator but why didn't we divide the tens?

OpenStudy (anonymous):

for example: (2 x10^2)/(4 x 10^3)= we divide the 2 and 4

OpenStudy (anonymous):

No, we don't.

OpenStudy (anonymous):

really the what do we do?

OpenStudy (anonymous):